Carlos Gonzales is production manager at an assembly plant that manufactures cordless telephones. The company is planning to install a pull system. The process is being planned to have a usage rate of 30 pieces per hour. Each container is designed to hold 10 pieces. It takes an average of 25 minutes to complete a cycle. How many containers will be needed? (Calculate to 2 decimal places)
Answer #1

Demand rate, D = 30 per hour

Container size, K = 10

Lead time, L = 25 minutes = 25/60 = 0.4167 hour

Number of containers needed = D*L/K

= 30*0.4167/10

= 1.25